OpenStudy (anonymous):

how do I convert x^2-y^2=4 to polar form?

OpenStudy (experimentx):

\[ x = r \cos \theta, y = r \sin\theta\]

OpenStudy (anonymous):

so how do I know what r and theta are?

OpenStudy (anonymous):

let x=rcostheta and y=rsin(thetha)

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=polar+plot+r^2+%3D+cos%282theta%29 R is the distance at angel theta from origin

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

then take out the common factor r^2, and then u knw that sin^2theta and cos^2theta =1

OpenStudy (anonymous):

oh yeah, cool guys thanks a lot i appreciate it

OpenStudy (anonymous):

ok cool man thank you